Mackenzie Company has a price oF $37 and will issue a dividend of $2.00 next year. It has a beta of 1.2​, the​ risk-free rate is 5.9%​, and the market risk premium is estimated to be 5.1%.

a. Estimate the equity cost of capital for Mackenzie.

b. Under the​ CGDM, at what rate do you need to expect​ Mackenzie's dividends to grow to get the same equity cost of capital as in part ​(a​)?
Answer #1

a. Equity cost of capital = Risk-free rate + Beta(Market risk premium)

Equity cost of capital = 0.059 + 1.2(0.051)

Equity cost of capital = 0.1202 or 12.02%

b. P0 = D1 / (r - g)

$37 = $2.00 / (0.1202 - g)

0.1202 - g = $2.00 / $37

0.1202 - g = 0.0541

g = 0.1202 - 0.0541

g = 0.0661 or 6.61%